# Tony Lindeberg

### v4

@@ 1 Introduction
Concerning vision, where one of the tasks is to recognize objects from the light that is reflected from them, it is clear that the identity of an object remains, even though we may view the object at different positions in the world, from different distances, different viewing directions and different relative motions between the object and the observer. Concerning the related task of recognizing spatio-temporal events, we may also regard different instantiations of spatio-temporal events, that occur either faster or slower relative to each other, as similar types of spatio-temporal events, under variations of the temporal scales in the time-varying phenomena. Concerning hearing, in a corresponding manner, it is clear that auditory stimuli may carry qualitatively similar information at different time moments, differing in phase and absolute pitch, or differing in the time frame over which the stimuli is presented. Given this basic property, that objects in the environment preserve their identity over classes of geometric transformations, one may hence ask if this property could be regarded as a determining principle for the evolution of visual perception.
@@
Compared to the previous publications on this topic, this presentation will be significantly more integrated, where the different types of variabilities that were addressed individually in in Lindeberg (2025c; 2025a) are here addressed in a unified manner, thus reframing the individually addressed variabilities in Lindeberg (2025c; 2025a) into a wider context. In relation to the early approach to such variabilities in Lindeberg (2023b), the resulting biological hypotheses will also be formulated based on significantly further theoretical depth, [^2] Beyond the first contributions in Lindeberg (2025c; 2025a), the theoretical background to the biological hypotheses in this paper will be based on the set of affine covariant derivative-based models for visual receptive fields to be described in Section 5.3.2, that were published after the publications of the earlier more biologically oriented contributions, and which will be built upon in the formulation of the more explicit affine-covariant models of visual receptive fields in Section 5.4 and 5.5, will then constitute a more solid foundation for the extended biological predictions to be presented in Section 6. In addition, we provide an alternative model for handling variabilities with respect to uniform spatial scaling transformations and temporal scaling transformations in Section 6, by only representing the finest spatial and temporal scales in the sensorium, as enabled by a formal theory for cascade smoothing properties, and as developed in the recent companion paper (Lindeberg 2025g). as enabled by the generalized mathematical theory presented in Lindeberg (2025d; 2025g).
@@
We will additionally describe a set of significant extensions [^3] The explicit statements in Sections 5.4, 5.5 and 6 go significantly further compared to the related results in the previous publications. relative to the previously presented theoretical results, as enabled by describing the overall theory in a unified manner relative to the previous more specialized technical contributions. Specifically, we will use results from this theory to address the topic of variabilities of simple cells [^4] According to the taxonomy of neurons in the primary visual cortex by Hubel and Wiesel (1959; 1962; 1968; 2005), a visual neuron is said to be simple if: (i) it has distinct excitatory and inhibitory subregions, (ii) it obeys roughly linear summation properties and (iii) the excitatory and inhibitory regions balance each other in diffuse lighting. in the primary visual cortex, based on a hypothesis that the shapes of the spatial or spatio-temporal receptive fields ought to span the degrees of the geometric image transformations that are involved in the image formation process, as well as provide further depth to the alternative working hypothesis that a vision system could choose to only implement the lowest levels of spatial and temporal scales in the sensorium corresponding to the simple cells. This will be done by making use of cascade smoothing properties over spatial and temporal scales, as enabled by the recently developed hybrid Lie semi-group theory over the filter shape parameters in the multi-parameter generalized Gaussian derivative model for visual receptive fields, presented in Lindeberg (2025g).
@@
Based on these theoretical results, we will then propose an extended set of theoretical predictions in Section 6, and relate them to the neurophysiological results that we currently find as available, and also propose new neurophysiological and psychophysical experiments to address the questions posed from this theoretical study, and which previous neurophysiological and psychophysical experiments may not have been designed to address.
@@ 1.1 Structure of this presentation
Then, Section 4 defines the axiomatically determined idealized models for visual receptive fields that we base this work upon, complemented in Section 5 by describing the covariance properties of the idealized receptive fields under the considered classes of geometric image transformations. Section 6 then addresses whether we can regard the spatial and the spatio-temporal shapes of simple cells in the primary visual cortex of higher mammals to span the variabilities of geometric image transformations, to support explicitly covariant families of visual receptive fields. Section 6 also describes ideas to future neurophysiological and psychophysical experiments to investigate this topic in more detail. Finally, Section 7 concludes with a summary and discussion.
@@ 2 Relations to previous work
Computational models of simple cells have specifically been expressed in terms of Gabor filters by Marcelja (1980), Jones and Palmer (1987a; 1987b), Porat and Zeevi (1988), Ringach (2002; 2004), Serre et al. (2007) and De and Horwitz (2021), and in terms of Gaussian derivatives by Koenderink and van Doorn (1984; 1987; 1992), Young (1987), Young et al. (2001; 2001) and Lindeberg (2013a; 2021b). Theoretical models of early visual processes have also been formulated based on Gaussian derivatives by Lowe (2000), May and Georgeson (2007), Hesse and Georgeson (2005), Georgeson et al. (2007), Hansen and Neumann (2008), Wallis and Georgeson (2009), Wang and Spratling (2016), Pei et al. (2016), Ghodrati et al. (2017), Kristensen and Sandberg (2021), Abballe and Asari (2022), Ruslim et al. (2023) and Wendt and Faul (2024).
@@
The main subject of this paper is to model the receptive fields of simple cells based on the normative theory for visual receptive fields proposed in Lindeberg (2021b) in terms of the generalized Gaussian derivative model, and then consider the influence on the resulting receptive field responses caused by variabilities in geometric image transformations.
@@ 3 Main classes of locally linearized geometric image transformations
Thereby, the composed image transformation model captures the variabilities of the scaled orthographic projection model, complemented with a variability over projections of 3-D motions between an observed object and the observer, including spatio-temporal events, that may occur faster or slower relative to a reference view, see Figure 3 for an illustration.
@@
Figure 6: Illustration of the variability of spatio-temporal receptive fields under Galilean transformations. Here, the mixed spatio-temporal derivative $T_{x{\bar{t}}}(x,t;\;s,\tau,v)=\partial_{x}\,\partial_{\bar{t}}\left(g(x-v\,t;\;s)\,\Psi(t;\;\tau,c)\right)$ over a 1+1-D spatio-temporal domain is shown for different values of the velocity parameter $v$, based on using a first-order Gaussian derivative over the spatial domain and a first-order derivative of the time-causal limit kernel over the temporal domain, for $s=\sigma_{x}^{2}$, $\tau=\sigma_{t}^{2}$ and $c=2$ with $\sigma_{x}=1/2$ and $\sigma_{t}=1$. This variability corresponds to varying the relative motion between the viewing direction and a moving local surface patch. (Horizontal axes: spatial coordinate $x\in[-3.5,3.5]$. Vertical axes: time $t\in[0,3.5]$.)
@@ 4.1 Receptive field models in terms of linear spatial or spatio-temporal convolution operations
A main subject of this paper is to present a set of theoretical extensions to this linear convolution structure, to make it possible for receptive fields with structurally similar properties as the simple cells in the primary visual cortex to handle more developed sets of geometric image transformations applied to the image data used as input to a vision system.
@@ 4.3 Idealized spatial or spatio-temporal models for simple cells in the primary visual cortex
The receptive fields according to this model have been obtained based on axiomatic derivations that reflect symmetry properties of the environment in combination with internal consistency requirements to guarantee theoretically well-founded treatment of image structures over different spatial and temporal scales. In this respect, the families of receptive fields have been formulated in a theoretically well-founded manner.
@@
where for the symbols not previously defined in connection with Equation (18) we have that:
@@
This form of the temporal smoothing function corresponds to using an infinite set of first-order integrators that are coupled in cascade, with the time constants chosen so as to specifically enable temporal scale covariance. The distribution parameter $c>1$, often chosen as $c=\sqrt{2}$ or $c=2$, in this temporal smoothing function reflects the ratio between adjacent discrete temporal scale levels in the corresponding temporal scale-space model.
In Lindeberg (2021b), it was demonstrated that idealized receptive field models of these types do rather well model the qualitative shape of biological simple cells as obtained by neurophysiological measurements by DeAngelis et al. (1995; 2004), Conway and Livingstone (2006) and Johnson et al. (2008); see Figures 12–18 in Lindeberg (2021b) for comparisons between biological receptive fields and idealized models thereof, based on the generalized Gaussian derivative model for visual receptive fields.
@@ 5.6 Relationships to the influence of illumination variations
Since the parameters $(S_{t},A,u,S_{t})$ of the geometric image transformation cannot be expected to be a priori known to a vision system that is to analyze an a priori unknown scene, a general purpose strategy for a vision system could therefore be to expand the receptive fields over a rich set of receptive fields, with the shapes of the receptive fields expanded over the degrees of freedom of the corresponding image transformations. Thereby, it would be possible to match the outputs from populations of receptive fields, to establish a corresponding matching of the receptive field responses obtained from a particular viewing condition in relation to a learned memory of receptive field responses computed from similar objects and spatio-temporal events under different sets of viewing conditions, see Figure 14 for an illustration. [^10] Note also that in an actual implementation of a matching scheme of receptive field responses of this form, one could also conceive that such a matching could be performed based on receptive field responses at higher levels in the visual hierarchy, which would indeed be possible based on the presented theory, if the receptive field responses at the higher levels are computed from the receptive field responses from the responses of the simple cells in a causal feed-forward manner that respects the covariance properties.
@@ 6.1 Purely spatial variabilities in the shapes of spatial and spatio-temporal receptive fields
Thereby, a vision system could in principle choose to compute the earliest layers of spatial receptive fields at only the finest spatial scale and nevertheless be able to compute the coarser spatial receptive fields at higher layers in the visual hierarchy. Thus, irrespective of whether the spatial receptive fields corresponding to the simple cells are expanded over the spatial scales, it seems very plausible that the vision system should have the ability to compute visual operations corresponding to spatial scale covariance.
@@ 6.2 Additional variabilities involving the temporal domain
Additionally, by extending the arguments in Lindeberg (2023b):
@@
From the ability of the simple cells in the visual system of higher mammals to compute spatio-temporal receptive fields similar to velocity-adapted temporal derivatives (DeAngelis et al. 1995; 2004; Lindeberg 2021b Figure 18 bottom part) and the receptive fields in the area MT being able to compute velocity-dependent responses, it seems plausible that the visual system should be able to compute Galilean-covariant receptive field responses.
Figure 6 in this paper shows an example of such a variability under Galilean transformations for spatio-temporal receptive fields over a 1+1-D spatio-temporal domain, based on a first-order spatial derivative of a Gaussian kernel and a first-order temporal derivative of the time-causal limit kernel.
@@
In the case of a non-causal temporal domain, the difference kernel $\Delta h$ is a Gaussian kernel $(\Delta h)(t;\;\tau_{1}\mapsto\tau_{2})=g(t;\;\tau_{2}-\tau_{1})$ because of the semi-group property of the Gaussian kernel. In the case of a time-causal domain, the difference kernel is either a single truncated exponential kernel $h_{\exp}(t;\;\mu_{k})=1/\mu_{k}\exp(-t/\mu_{k})$ if $t>0$, or 0 otherwise, or the convolution of a set of such truncated exponential kernels, with appropriate time constants $\mu_{k}$ to fit the relationship between the temporal scale levels $\tau_{1}$ and $\tau_{2}$.
@@
An extended theory for how an idealized vision system, based on receptive fields in terms of the generalized Gaussian derivative theory for visual receptive fields, could compute receptive field responses at coarser spatial and temporal scales, based on cascade smoothing properties from finer spatial and temporal scales, is given in Lindeberg (2025g).
@@
    For efficiency reasons, separate expansions over the affine and the Galilean degrees of freedom would need much less wetware to span the degrees of freedom of the corresponding geometric image transformations. A joint expansion over these degrees of freedom could on the other hand allow for the computation of more accurate estimates of the structure of dynamic scenes.
@@
  • How wide ranges in the parameter spaces of the receptive fields are spanned for the respective degrees of freedom for different species?
Answering these questions could provide valuable cues about how well different types of species have adapted their visual systems to handle the influence of different types of geometric image transformations on the image data.
@@
Currently, it seems very hard to get access to raw reconstructed receptive field profiles from previous recordings of the visual receptive fields in V1 of higher mammals. Publicly available receptive fields, as published in scientific publications, typically only show a rather small number of visual receptive fields. Concerning the hypothesized variabilities with regard to the degree of elongation of visual receptive fields, the neurophysiological evidence, that we have based our previous hypothesis on, are based on orientation selectivity properties of higher mammals (primates or cats) with a clear pinwheel structure. Recordings from mice, for which there is the largest number of public data, may on the other hand not be able to fully answer this question, because the vision system of mice is far less developed compared the visual systems of higher mammals (see e.g. Huberman and Niell (2011)), and does specifically lack pinwheel structures. In addition, recent results by Fu et al. (2024) of reconstructing most exciting inputs (MEIs) for early visual receptive fields in mice suggest that complex spatial features emerge earlier in the visual pathway of mice compared to primates, and that the receptive fields of mice are therefore more complex than a Gabor or Gaussian derivative model, as used in the theoretical model for stating the predictions in this paper.
For these overall reasons, in Lindeberg (2023b) Sections 3.2.1–3.2.2 and Lindeberg (2025c) Sections 4.2–4.3, sets of suggestions for more specific neurophysiological experiments are proposed to investigate the underlying hypotheses in more detail, and to characterize the structure of biological receptive fields in the primary visual cortex with respect to the influence of parameters of the receptive fields corresponding to the different degrees of freedom of the 4 main types of geometric image transformations. Complementary suggestions for further research in relation to the influence of geometric image transformations on early vision are also outlined in Lindeberg (2025a) Section 8.1.
Additionally, in Lindeberg (2025e) it is shown that that the results concerning an expansion over the degree of elongation for simple cells appear to extend to complex cells.

### v5

@@ 1 Introduction
Concerning vision, where one of the tasks is to recognize objects from the light that is reflected from them, it is clear that the identity of an object remains, even though we may view the object at different positions in the world, from different distances, different viewing directions and different relative motions between the object and the observer. Concerning the related task of recognizing spatio-temporal events, we may also regard different instantiations of spatio-temporal events, that occur either faster or slower relative to each other, as similar types of spatio-temporal events, under variations of the temporal scales in the time-varying phenomena. Concerning hearing, in a corresponding manner, it is clear that auditory stimuli may carry qualitatively similar information at different time moments, differing in absolute pitch and phase, or differing in the time frame over which the stimuli is presented. Given this basic property, that objects in the environment preserve their identity over classes of geometric transformations, one may hence ask if this property could be regarded as a determining principle for the evolution of visual perception.
@@
Compared to the previous publications on this topic, this presentation will be significantly more integrated, where the different types of variabilities, that were addressed individually in in Lindeberg (2025c; 2025a), are here addressed in a unified manner, thus reframing the individually addressed variabilities in Lindeberg (2025c; 2025a) into a wider context. In relation to the early approach to such variabilities in Lindeberg (2023b), the resulting biological hypotheses will also be formulated based on significantly further theoretical depth, [^2] Beyond the first contributions in Lindeberg (2025c; 2025a), the theoretical background to the biological hypotheses in this paper will be based on the set of affine covariant derivative-based models for visual receptive fields to be described in Section 5.3.2, that were published after the publications of the earlier more biologically oriented contributions, and which will be built upon in the formulation of the more explicit affine-covariant models of visual receptive fields in Section 5.4 and 5.5, will then constitute a more solid foundation for the biological predictions to be presented in Section 6, with specifically substantial extensions regarding the predictions about variabilities with respect to Galilean transformations and temporal scaling in Section 6.2 and the formulation of overall neuroscientific questions in Section 6.3. In addition, we provide an alternative model for handling variabilities with respect to uniform spatial scaling transformations and temporal scaling transformations in Section 6, by only representing the finest spatial and temporal scales in the sensorium, as enabled by a formal theory for cascade smoothing properties, and as developed in the recent companion paper (Lindeberg 2025h). as enabled by the generalized mathematical theory presented in Lindeberg (2025d; 2025h) and the complementary analysis of the directional selectivity properties of the idealized models of the spatio-temporal receptive fields of simple cells recently performed in Lindeberg (2025g).
@@
We will additionally describe a set of significant extensions [^3] The explicit statements in Sections 5.4, 5.5 and 6 go significantly further compared to the related results in the previous publications. relative to the previously presented theoretical results, as enabled by describing the overall theory in a unified manner relative to the previous more specialized technical contributions. Specifically, we will use results from this theory to address the topic of variabilities of simple cells [^4] According to the taxonomy of neurons in the primary visual cortex by Hubel and Wiesel (1959; 1962; 1968; 2005), a visual neuron is said to be simple if: (i) it has distinct excitatory and inhibitory subregions, (ii) it obeys roughly linear summation properties and (iii) the excitatory and inhibitory regions balance each other in diffuse lighting. in the primary visual cortex, based on a hypothesis that the shapes of the spatial or spatio-temporal receptive fields ought to span the degrees of the geometric image transformations that are involved in the image formation process, as well as provide further depth to the alternative working hypothesis that a vision system could choose to only implement the lowest levels of spatial and temporal scales in the sensorium corresponding to the simple cells. This will be done by making use of cascade smoothing properties over spatial and temporal scales, as enabled by the recently developed hybrid Lie semi-group theory over the filter shape parameters in the multi-parameter generalized Gaussian derivative model for visual receptive fields, presented in Lindeberg (2025h).
@@
Based on these theoretical results, we will then propose an extended set of theoretical predictions in Section 6, and relate them to the neurophysiological results, that we currently find as available, and also propose new neurophysiological and psychophysical experiments to address the questions posed from this theoretical study, and which previous neurophysiological and psychophysical experiments may not have been designed to address.
@@ 1.1 Structure of this presentation
Then, Section 4 defines the axiomatically determined idealized models for visual receptive fields, that we base this work upon, complemented in Section 5 by describing the covariance properties of the idealized receptive fields under the considered classes of geometric image transformations. Section 6 then addresses whether we can regard the spatial and the spatio-temporal shapes of simple cells in the primary visual cortex of higher mammals to span the variabilities of geometric image transformations, to support explicitly covariant families of visual receptive fields. Section 6 also describes ideas to future neurophysiological and psychophysical experiments to investigate this topic in more detail. Finally, Section 7 concludes with a summary and discussion.
@@ 2 Relations to previous work
Computational models of simple cells have specifically been expressed in terms of Gabor filters by Marcelja (1980), Jones and Palmer (1987a; 1987b), Porat and Zeevi (1988), Ringach (2002; 2004), Serre et al. (2007), Baspinar et al. (2018; 2020) and De and Horwitz (2021), and in terms of Gaussian derivatives by Koenderink and van Doorn (1984; 1987; 1992), Young (1987), Young et al. (2001; 2001) and Lindeberg (2013a; 2021b). Theoretical models of early visual processes have also been formulated based on Gaussian derivatives by Lowe (2000), May and Georgeson (2007), Hesse and Georgeson (2005), Georgeson et al. (2007), Hansen and Neumann (2008), Wallis and Georgeson (2009), Wang and Spratling (2016), Pei et al. (2016), Ghodrati et al. (2017), Kristensen and Sandberg (2021), Abballe and Asari (2022), Ruslim et al. (2023) and Wendt and Faul (2024).
@@
The main subject of this paper is to model the receptive fields of simple cells based on the normative theory for visual receptive fields proposed in Lindeberg (2021b), in terms of the generalized Gaussian derivative model, and then consider the influence on the resulting receptive field responses caused by variabilities in geometric image transformations.
@@ 3 Main classes of locally linearized geometric image transformations
Thereby, the composed image transformation model captures the variabilities of the scaled orthographic projection model, complemented with a variability over projections of 3-D motions between an observed object and the observer, including spatio-temporal events, that may occur faster or slower relative to a previously observed reference view, see Figure 3 for an illustration.
@@
Figure 6: Illustration of the variability of spatio-temporal receptive fields under Galilean transformations. Here, the second-order spatial derivative $T_{xx}(x,t;\;s,\tau,v)=\partial_{x}\,\partial_{\bar{t}}\left(g(x-v\,t;\;s)\,\Psi(t;\;\tau,c)\right)$ over a 1+1-D spatio-temporal domain is shown for different values of the velocity parameter $v$, based on using a first-order Gaussian derivative over the spatial domain and the zero-order derivative of the time-causal limit kernel over the temporal domain, for $s=\sigma_{x}^{2}$, $\tau=\sigma_{t}^{2}$ and $c=2$ with $\sigma_{x}=1$ and $\sigma_{t}=1$. This variability corresponds to varying the relative motion between the viewing direction and a moving local surface patch. (Horizontal axes: spatial coordinate $x\in[-5,5]$. Vertical axes: time $t\in[0,5]$.)
@@ 4.1 Receptive field models in terms of linear spatial or spatio-temporal convolution operations
A main subject of this paper is to present a set of theoretical extensions to this linear convolution structure, to make it possible for receptive fields with structurally similar properties as the simple cells in the primary visual cortex, to handle more developed sets of geometric image transformations applied to the image data used as input to a vision system.
@@ 4.3 Idealized spatial or spatio-temporal models for simple cells in the primary visual cortex
The receptive fields according to this model have been obtained based on axiomatic derivations, that reflect symmetry properties of the environment in combination with internal consistency requirements to guarantee theoretically well-founded treatment of image structures over different spatial and temporal scales. In this respect, the families of receptive fields have been formulated in a theoretically well-founded manner.
@@
where, for the symbols not previously defined in connection with Equation (18), we have that:
@@
This form of the temporal smoothing function corresponds to using an infinite set of first-order integrators that are coupled in cascade, with the time constants chosen so as to specifically enable temporal scale covariance. The distribution parameter $c>1$, often chosen as $c=\sqrt{2}$ or $c=2$, in this temporal smoothing function reflects the ratio between adjacent discrete temporal scale levels in the corresponding temporal scale-space model. (For practical implementation purposes, the infinite number of first-order integrators can often be truncated after the 4 to 8 slowest time constants.)
In Lindeberg (2021b), it was demonstrated that idealized receptive field models of these types do rather well model the qualitative shape of biological simple cells, as obtained by neurophysiological measurements by DeAngelis et al. (1995; 2004), Conway and Livingstone (2006) and Johnson et al. (2008); see Figures 12–18 in Lindeberg (2021b) for comparisons between biological receptive fields and idealized models thereof, based on the generalized Gaussian derivative model for visual receptive fields.
@@ 5.6 Relationships to the influence of illumination variations
Since the parameters $(S_{t},A,u,S_{t})$ of the geometric image transformation cannot be expected to be a priori known to a vision system, that is to analyze an a priori unknown scene, a general purpose strategy for a vision system could therefore be to expand the receptive field measurements over a rich family of receptive fields, with the shapes of the receptive fields expanded over the degrees of freedom of the corresponding image transformations. Thereby, it would be possible to match the outputs from populations of receptive fields, to establish a corresponding matching of the receptive field responses obtained from a particular viewing condition in relation to a learned memory of receptive field responses computed from similar objects and spatio-temporal events under different sets of viewing conditions, see Figure 14 for an illustration. [^10] Note also that in an actual implementation of a matching scheme of receptive field responses of this form, one could also conceive that such a matching could be performed based on receptive field responses at higher levels in the visual hierarchy, which would indeed be possible based on the presented theory, if the receptive field responses at the higher levels are computed from the receptive field responses from the responses of the simple cells in a causal feed-forward manner that respects the covariance properties.
@@ 6.1 Purely spatial variabilities in the shapes of spatial and spatio-temporal receptive fields
Thereby, a vision system could in principle choose to compute the earliest layers of spatial receptive fields at only the finest spatial scale and nevertheless be able to compute the coarser spatial receptive fields at higher layers in the visual hierarchy. [^11] For example, regarding variabilities in receptive field size in higher layers of the visual hierarchy, Orban (1997) (page 371) reports that in the visual area MT, “… at all eccentricities there is a tenfold range in RF size”. Thus, irrespective of whether the spatial receptive fields corresponding to the simple cells are expanded over the spatial scales, it seems very plausible that the vision system should have the ability to compute visual operations corresponding to spatial scale covariance.
@@
Spatial differentiation order $m=1$
$\theta=0$ $\theta=\pi/3$ $\theta=2\pi/3$
Refer to caption Refer to caption Refer to caption
$\theta=\pi$ $\theta=4\pi/3$ $\theta=5\pi/3$
Refer to caption Refer to caption Refer to caption
Spatial differentiation order $m=2$
$\theta=0$ $\theta=\pi/3$ $\theta=2\pi/3$
Refer to caption Refer to caption Refer to caption
$\theta=\pi$ $\theta=4\pi/3$ $\theta=5\pi/3$
Refer to caption Refer to caption Refer to caption
Figure 18: Variability in the orientation selectivity properties of idealized models of simple cells of the form (4.3) in terms of velocity-adapted affine Gaussian kernels for spatial derivative orders $m=1$ and $m=2$ and temporal derivative order $n=0$ over the direction $\theta=\arg v$ of the image velocity $v$ for a set of velocity-adapted spatio-temporal receptive fields in the directions $\theta\in\{0,\pi/3,2\pi/3,\pi,4\pi/3,5\pi/3\}$. The different colours of the orientation selectivity curves show the results for different degrees of elongation $\kappa\in\{1,2,4,8\}$, with the directional selectivity curves becomes sharper for larger values of $\kappa$. All the orientation selectivity curves have been computed using the same magnitude $|v|$ of the image velocity, only by varying the direction $\theta=\arg v$. The magnitudes of each directional selectivity curve have, in turn, been normalized such that they assume the maximum value 11 when velocity of the stimulus is equal to the velocity parameter of the receptive field. For other velocity magnitude values of the stimulus, the peak value of the directional selectivity curve will be lower.
@@ 6.2 Additional variabilities involving the temporal domain
Additionally, by extending the arguments in Lindeberg (2023b), we can predict the following variabilities with regard to time-dependent image data:
@@
Regarding processing of time-dependent image data in the primary visual cortex and the middle temporal visual area MT, we have the following facts (Lindeberg (2025g):
  • the ability of the simple cells in the visual system of higher mammals to compute spatio-temporal receptive field responses similar to velocity-adapted temporal derivatives (DeAngelis et al. 1995; 2004; de Valois et al. 2000; Lindeberg 2021b Figure 18 bottom part),
  • visual neurons in the primate visual cortex being known to be direction selective (Hubel 1959, Orban et al. 1986, Churchland et al. 2005), and “… organized into subcolumns within an iso-orientation column, with each subcolumn preferring motion in a different direction” (Elstrott and Feller 2009),
  • the receptive fields in the middle temporal visual area MT being able to compute to a rich variety of direction-selective responses (Orban 1997, Born and Bradley 2005),
  • with the output from the primary visual cortex providing input to the middle temporal visual area MT (Movshon and Newsome 1996) and
  • with the direction-selective neurons in MT organized into direction columns (Albright et al. 1984).
From these results, it appears as if we can regard the primary visual cortex and middle temporal visual area MT as performing expansions of the visual representations over the directions and speeds of local Galilean motions. Thereby, it seems plausible that:
  • the visual system should be able to compute Galilean-covariant receptive field responses, so as to be able to process visual information over a wide range of motion speeds [^12] The range of velocities may, however, be strongly dependent on the distance from the center of the fovea, as characterized in detail by Orban et al. 1986, and probably depending on both the receptive field size, which roughly increases linearily from the distance from the center of the fovea, and the minimum temporal scale of the neuron. and motion directions,
  • this motion processing hierarchy could then be based on Galilean-covariant receptive fields in the primary visual cortex.
Figure 6 in this paper shows an example of such a variability under Galilean transformations for spatio-temporal receptive fields over a 1+1-D spatio-temporal domain, based on a first-order spatial derivative of a Gaussian kernel and a first-order temporal derivative of the time-causal limit kernel. Figure 18 shows the variability in the directional selectivity properties of simple cells that such a variability in the image velocity $v$ of the receptive fields gives rise to, when also including a variability over the degree of elongation $\kappa$ of the receptive fields, as based on the theoretical analysis in Lindeberg (2025g).
@@
In the case of a non-causal temporal domain, the difference kernel $\Delta h$ is a Gaussian kernel $(\Delta h)(t;\;\tau_{1}\mapsto\tau_{2})=g(t;\;\tau_{2}-\tau_{1})$ because of the semi-group property of the Gaussian kernel. In the case of a time-causal domain, the difference kernel is either a single truncated exponential kernel $h_{\exp}(t;\;\mu_{k})=1/\mu_{k}\exp(-t/\mu_{k})$ if $t\geq 0$, or 0 otherwise, or the convolution of a set of such truncated exponential kernels, with appropriate time constants $\mu_{k}$ to fit the relationship between the temporal scale levels $\tau_{1}$ and $\tau_{2}$, according to the recurrence relation in Lindeberg (2023a) Equation (28).
@@
An extended theory for how an idealized vision system, based on receptive fields in terms of the generalized Gaussian derivative theory for visual receptive fields, could compute receptive field responses at coarser spatial and temporal scales, based on cascade smoothing properties from finer spatial and temporal scales, is given in Lindeberg (2025h).
@@
    For efficiency reasons, separate expansions over the affine and the Galilean degrees of freedom would need much less wetware to span the degrees of freedom of the corresponding geometric image transformations. A joint expansion over these degrees of freedom could on the other hand allow for the computation of more accurate estimates of the 3-D structure of dynamic scenes.
@@
  • For which combinations of spatial and temporal derivative operations according to the theory for generalized Gaussian derivative based spatial and spatio-temporal receptive fields can biological receptive fields be found with qualitatively similar shapes?
  • How wide ranges in the parameter spaces of the receptive fields are spanned for the respective degrees of freedom for different species? Does the vision system span dense sets over the multi-dimensional parameter space, or does the vision system only span sparser submanifolds, with possibly coupled dependencies between different subsets of the parameters?
Answering these questions could provide valuable cues about how well different types of species have adapted and organized their visual systems to handle the influence of different types of geometric image transformations on the image data.
@@
Currently, it seems very hard to get access to raw reconstructed receptive field profiles from previous recordings of the visual receptive fields in V1 of higher mammals. Publicly available receptive fields, as published in scientific publications, typically only show a rather small number of visual receptive fields. Concerning the hypothesized variabilities with regard to the degree of elongation of visual receptive fields, the neurophysiological evidence, that we have based our previous hypothesis on, are based on orientation selectivity properties of higher mammals (primates or cats) with a clear pinwheel structure. Our support for Galilean covariance is correspondingly largely based on measurements of direction and speed selecitivity properties of motion-sensitive neurons.
Recordings from mice, for which there is the largest number of public data, may on the other hand not be able to fully answer this question, because the vision system of mice is far less developed compared the visual systems of higher mammals (see e.g. Huberman and Niell (2011)), and does specifically lack pinwheel structures. In addition, recent results by Fu et al. (2024) of reconstructing most exciting inputs (MEIs) for early visual receptive fields in mice suggest that complex spatial features emerge earlier in the visual pathway of mice compared to primates, and that the receptive fields of mice are therefore more complex than a Gabor or Gaussian derivative model, as used in the theoretical model for stating the predictions in this paper.
For these overall reasons, in Lindeberg (2023b) Sections 3.2.1–3.2.2, Lindeberg (2025c) Sections 4.2–4.3 and Lindeberg (2025g) Section 8, sets of suggestions for more specific neurophysiological experiments are proposed to investigate the underlying hypotheses in more detail, and to characterize the structure of biological receptive fields in the primary visual cortex with respect to the influence of parameters of the receptive fields corresponding to the different degrees of freedom of the 4 main types of geometric image transformations. Complementary suggestions for further research in relation to the influence of geometric image transformations on early vision are also outlined in Lindeberg (2025a) Section 8.1.
Additionally, concerning the extension to non-linear visual neurons, in Lindeberg (2025e) it is shown that that the results concerning an expansion over the degree of elongation for simple cells appear to extend to complex cells.